OpenStudy (anonymous):

The slope of a line is -4, and the y-intercept is -3. What is the equation of the line written in slope-intercept form? y = -4x - 3 y = -4x + 3 y = 4x - 3

OpenStudy (ajprincess):

slope intercept form of equation is y=mx+c m-slope c-y-intercept so what is m and c?

OpenStudy (anonymous):

idk im so tired right now i cant think

OpenStudy (ajprincess):

what is the slope given?

OpenStudy (anonymous):

-4

OpenStudy (ajprincess):

ya good:) nw that is ur m. what is the y-intercept?

OpenStudy (anonymous):

-3

OpenStudy (ajprincess):

ya good:) nw that is ur c.

OpenStudy (anonymous):

true statement is y=-4x-3

OpenStudy (ajprincess):

plug in those values of m and c in y=mx+c nd u will get the required answer

OpenStudy (anonymous):

y=-4 -3

OpenStudy (ajprincess):

ya it is y=-4x-3

OpenStudy (anonymous):

slope m=-4,y intercept c=-3 eq. of line in intercept form is y=mx+c eq. is y=-4x-3